[Hallc_running] [New Logentry] RC Update

whit at jlab.org whit at jlab.org
Thu Oct 25 21:25:03 EDT 2018

Logentry Text:
* We started taking data around 5:00 pm. 
* After looking at the harp scans and Yag Viewer, the beam is good for production.
* Note this problem will need attention https://logbooks.jlab.org/entry/3614867
* We will  continue production through the owl shift.
* Currently we are on Kinematic 13-3.
* The rates in SHMS at 14 degrees (33uA instead of 58 uA to get 500 kHz). At 16 degrees we ran at 60uA to get 500kHz.
Shift workers are still needed! Please sign up. Please clap.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3614948
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181025/89e4acd6/attachment.html>

More information about the Hallc_running mailing list